OpenStudy (anonymous):

Trying to simplify: (Square root x-2) + 5 divided by (square root x-2) all over x-2. Can anyone help? Thanks!

OpenStudy (anonymous):

I hate these.

OpenStudy (anonymous):

sqrt(x-2)+5 / sqrt(x-2)(x-2)....yes?

OpenStudy (anonymous):

Multiplying top and bottom by sqrt(x-2) looks like a good bet...

OpenStudy (anonymous):

No, ((\sqrt{x-2})+ 5/\sqrt{x-2}) / x-2\]

OpenStudy (anonymous):

That's the same...

OpenStudy (anonymous):

1/2 divided by 2 is 1/4...

OpenStudy (anonymous):

\[\sqrt{x-2}+5 \over \sqrt{x-2}(x-2)\]

OpenStudy (anonymous):

sorry its backwards but thats what im trying to solve

OpenStudy (anonymous):

Are they all 2's in the picture?

OpenStudy (anonymous):

except for the 5 on the very top

OpenStudy (anonymous):

Well, that's straightforward, just add 'em 6 sqrt(x-2) /5(x-2)

OpenStudy (anonymous):

U see?

OpenStudy (anonymous):

maybe? - working on where you got the 6

OpenStudy (anonymous):

Just add at the top, LCD is 5 so u have 5 +1 sqrt(x-2)'s

OpenStudy (anonymous):

and the 5 goes to the bottom along with (x-29

OpenStudy (anonymous):

I see. Thanks.

OpenStudy (anonymous):

x-2, sorry

OpenStudy (anonymous):

Wait a minute I am reading it upside down....

OpenStudy (anonymous):

It actually says sqrt(x-2) + 5/sqrt(x-2) is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ah Ok, moment..

OpenStudy (anonymous):

OK; LCD is sqrt(x-2) so adding (x+3) / (x-2) sqrt(x-2) is what u want to simplify....

OpenStudy (anonymous):

So multiply top and bottom by sqrt(x-2) to get sqrt(x-2) (x+3) / (x-2)^2 is your answer.

OpenStudy (anonymous):

Where does the (x+3) come from?

OpenStudy (anonymous):

Your LCD is sqrt(x-2) so when you add you get sqrt(x-2)sqrt(x-2) +5 = x-2 +5 = x+3

OpenStudy (anonymous):

i finally see

OpenStudy (anonymous):

:-) Got there in the end (try to use the equation editor, makes life easier with the odd ones).

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

estudier's solution boils down to\[\frac{x+3}{(x-2)^{3/2}} \]if the numerator and the denominator are divided by\[\sqrt{(x-2) } \]